what pressure will be exerted by o.450 mol of a gas at 25 degree celcuis if it is contained in a vessel whose volume is 0.650
1
Expert's answer
2017-11-01T14:14:07-0400
Volume of the gas at 1 atm and 25°C: V1 = 0.450 mol ∙ 22.4 l/mol = 10.08 l
Volume of the vessel: V2 = 0.650 l
p ∙ V = const, therefore:
Pressure: (V1 / V2) ∙ 1 atm = (10.08 / 0.650) ∙ 1 atm = 15.5 atm